Determine whether a triangle can be formed with the given side lengths. If so, use Heron's formula to find the area of the triangle.
a = 240
b = 123
c = 207

Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

Each length is less than the sum of the other two sides and greater than the difference of the two other sides.
Yes, this can make a triangle.
Area is sqrt (s(s-a)(s-b)(s-c) where s=(1/2)(a+b+c)=(1/2)(570)=285
A=sqrt(285*45*162*78)=12,730.1 sq units.

Answer by ikleyn(52925)

.
On Heron's formula see the lessons
    - Proof of the Heron's formula for the area of a triangle,
    - One more proof of the Heron's formula for the area of a triangle,
in this site.


Also,  you have this free of charge online textbook on Geometry
    GEOMETRY - YOUR ONLINE TEXTBOOK
in this site.

The referred lessons are the part of this online textbook under the topic  "Area of triangles ".